# Topology of Discrete Quantum Feedback Control

### v2

@@ VII.3 QND demon
Here, we present a model of topological feedback control with strong unitary symmetry by following the construction in Sec. VI.3. We consider a spin-$1/2$ particle on a 1D lattice with two sublattices. A quantum state at site $j$ with sublattice $a=A,B$ and spin $\sigma=\uparrow,\downarrow$ is denoted by |$\ket{j,a,\sigma}$. We perform a projective position measurement on this particle, and assume that this measurement does not disturb the spin state of the particle. Thus, a measurement outcome is specified only by the site index $j$ and the sublattice index $a$, and therefore the measurement operator is given by
@@
is the magnetization. Thus, the position measurement does not influence the spin degrees of freedom and serves as a QND measurement for $\hat{S}^{z}$ (see Sec. VI.3.1).
@@
where $J_{1},J_{2}$ are the hopping amplitudes, $V$ is the height of the potential, and $h_{z}$ ($h_{s}$) denotes the strength of a uniform (staggered) magnetic field. The feedback Hamiltonian conserves the magnetization, i.e., $[\hat{H}_{m},\hat{S}^{z}]=0$. The Hamiltonians (146) and (147) are generalized versions of the chiral Maxwell’s demon discussed in Sec. IV. The CPTP map for this feedback control is given by Eq. (5) with $\hat{K}_{m}=\hat{U}_{m}\hat{M}_{m}$ and $\hat{U}_{m}=\exp[-i\hat{H}_{m}\tau]$.
Since $[\hat{M}_{m},\hat{S}^{z}]=[\hat{H}_{m},\hat{S}^{z}]=0$, the CPTP map for this feedback control has strong symmetry (see Sec. VI.3). Hence, the CPTP map can be decomposed into four parts which act on subspaces with different symmetry eigenvalues. Specifically, we have
@@
In Fig. 16, we plot the eigenspectra of the quantum channel $\tilde{\mathcal{E}}$ under the PBC and the OBC together with those of ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$. The topological characterization of the quantum channel can be applied to each symmetry sector. The PBC eigenspectrum of ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$ with $\sigma=\sigma^{\prime}$ contains a single loop with winding number $-2$, while that of ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$ with $\sigma\neq\sigma^{\prime}$ contains two loops, each of which has winding number $-1$. Here, the uniform and staggered magnetic fields lift the degeneracy of the eigenspectra of ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$’s. Importantly, the eigenspactrum of ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$ with $\sigma\neq\sigma^{\prime}$ is no longer symmetric with respect to the real axis, implying that the modular conjugation symmetry is broken in these symmetry sectors. Thus, ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$ with $\sigma\neq\sigma^{\prime}$ offers a topologically nontrivial superoperator in class A.
Figure 16: Eigenspectrum of a quantum channel for feedback control with strong unitary symmetry. In each figure, blue curves and the blue dot at the origin constitute the eigenspectrum under the PBC in the infinite-size limit, and orange dots form the eigenspectrum under the OBC with the number of unit cells $L=40$. (a) Full eigenspectrum of the CPTP map $\tilde{\mathcal{E}}$. (b) Eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\uparrow}$, which is identical to that of $\tilde{\mathcal{E}}_{\downarrow\downarrow}$. (c) Eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\downarrow}$. (d) Eigenspectrum of $\tilde{\mathcal{E}}_{\downarrow\uparrow}$. The paramaters of the model are set to $J_{1}=J_{2}=1$, $V=+\infty$, $h_{z}=0.5$, and $h_{s}=1$. The duration of feedback is $\tau=2$.
@@ VII.4 $\mathbb{Z}_{2}$ demon
As shown in Sec. VI.3.3, a quantum channel in class AII requires unitary symmetry and $C$ symmetry with $\eta_{C}=-1$ of the truncated Bloch matrix. To construct feedback control that fulfills these requirements, we consider a spin-$1/2$ particle on a 1D lattice with two sublattices as in Sec. VII.3. We first perform a spin-nondisturbing projective measurement of the position and the sublattice, where the measurement operator is given by Eq. (144). For a measurement outcome $m_{1}=(j_{1},a_{1})$, we perform feedback control with Hamiltonian $\hat{H}_{m_{1}}$ that has a symmetry
@@
An example of such feedback Hamiltonians is given by
@@
which are schematically illustrated in Fig. 17. For computational simplicity, here we assume that $\hat{H}_{j_{1},a}$ acts only on three unit cells labeled by $j_{1}-1,j_{1},j+1$; however, this assumption is not essential. After the unitary time evolution described by $\hat{U}_{j_{1},a_{1}}=\exp[-i\hat{H}_{j_{1},a_{1}}\tau]$, we perform the second projective measurement of the position and the sublattice with the same measurement operators (144) and suppose that we obtain an outcome $m_{2}=(j_{2},a_{2})$. Then, using the set of measurement outcomes $m_{1}$ and $m_{2}$, we perform additional feedback control with the following unitary operator:
$\hat{U}_{m_{1},m_{2}}^{\prime}=\begin{cases}\exp[-i\pi\hat{S}^{z}]&\mathrm{if}\ j_{1}>j_{2}\ \mathrm{and}\ a_{1}\neq a_{2};\\ \hat{I}&\mathrm{otherwise}.\end{cases}$ (155)
This addtional feedback control can be realized by a magnetic field and imprints a phase depending on the spin degrees of freedom. This feedback control is essential for realizing $C$ symmetry of a quantum channel (see below).
Figure 17: Schematic illustration of the $\mathbb{Z}_{2}$ demon. We first perform a projective measurement of the position and the sublattice without disturbing the spin degrees of freedom. (a) If a particle is detected at site $j_{1}$ and sublattice $A$, we perform feedback control by using the Hamiltonian $H_{j_{1},A}$. (b) If a particle is detected at site $j_{1}$ and sublattice $B$, we perform feedback control by using the Hamiltonian $H_{j_{1},B}$. After the unitary evolution under the feedback Hamiltonian $H_{j_{1},a_{1}}$ $(a_{1}=A,B)$, we perform the second projective measurement and additional feedback with a magnetic field [see Eq. (155)].
@@
The unitary operator with the feedback Hamiltonian $\hat{H}_{m_{1}}$ statisfies the following reciprocity relations because of the symmetry between $\hat{H}_{j_{1},A}$ and $\hat{H}_{j_{1},B}$:
@@
which follows from the symmetry (151) of the feedback Hamiltonian. Then, conditions (111a)-(111c) are satisfied, and thus ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$ belongs to class AII for $\sigma\neq\sigma^{\prime}$ and class AI $+$ psH+ for $\sigma=\sigma^{\prime}$.
The topological feedback control in class AII exhibits the $\mathbb{Z}_{2}$ skin effect [72]. To see this, we show the eigenspectrum of the quantum channel for the $\mathbb{Z}_{2}$ demon under the PBC and the OBC in Fig. 18. The block-diagonalized CPTP map $\tilde{\mathcal{E}}_{\uparrow\uparrow}$ does not possess $C$ symmetry with $\eta_{C}=-1$, and therefore the Kramers degeneracy of eigenvalues is not ensured. Consequently, the loop structure in the eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\uparrow}$ shown in Fig. 18(a) collapses near the real axis due to real-complex transitions and does not support nontrivial topology. In contrast, the block $\tilde{\mathcal{E}}_{\uparrow\downarrow}$ has $C$ symmetry with $\eta_{C}=-1$, leading to the Kramers-degenerated eigenspectrum shown in Fig. 18(b). The point-gap topology of this spectrum is characterized by a $\mathbb{Z}_{2}$ topological invariant [72]
@@
Indeed, the eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\downarrow}$ under the OBC forms a line rather than a closed loop, which is a hallmark of the $\mathbb{Z}_{2}$ skin effect. Interestingly, the $\mathbb{Z}_{2}$ skin effect can occur only in ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$ with $\sigma\neq\sigma^{\prime}$ since ~$\tilde{\mathcal{E}}_{\sigma\sigma}$ has the modular conjugation symmetry and is therfore characterized by the winding number rather than the $\mathbb{Z}_{2}$ invariant (see Sec. VII.2). Physically, the eigenspectra in Fig. 18 indicate that off-diagonal elements of the density matrix in terms of the $\hat{S}^{z}$ basis are accumulated on the edges of the system, whereas diagonal elements are delocalized over the system. Such distinct behavior is induced by the additional feedback in Eq. (155), which nontrivially acts only on the off-diagonal elements.
Figure 18: Eigenspectrum of a quantum channel for the $\mathbb{Z}_{2}$ demon. In each figure, blue curves and the blue dot at the origin constitute the eigenspectrum under the PBC in the infinite-size limit, and orange dots form the eigenspectrum under the OBC with the number of unit cells $L=20$. (a) Eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\uparrow}$. (b) Eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\downarrow}$. The paramaters of the model are set to $J_{1}=0.8,J_{2}=1$ and, $\lambda=0.5$. The duration of feedback is $\tau=1$.
Here we remark on the boundary condition of the $\mathbb{Z}_{2}$ demon. To see the non-Hermitian skin effect in symmetry-protected topological feedback control, a quantum channel under the OBC must respect the protecting symmetry. We detail how to implement a symmetry-preserving OBC in Appendix E.
@@
In Fig. 19, we plot the number density and the transeverse spin components of a particle after $N$ repeated applications of the feedback control with the $\mathbb{Z}_{2}$ demon. The numerical results clearly show that the transverse spin components are localized at the edges due to the skin effect, whereas the number density is delocalized over the system. Thus, the $\mathbb{Z}_{2}$ skin effect of the topological feedback control manifests itself only in the off-diagonal components of the density matrix. Note that the transverse spin components are not conserved due to the additional feedback in Eq. (155) and vanish in the steady state of the quantum channel [see Fig. 19 (g) and (h)]. Therefore, the nontrivial topology of the off-diagonal sectors appear only in the transient dynamics under repeated applications of the feedback control.
Figure 19: Dynamical signatures of nontrivial topology of a $\mathbb{Z}_{2}$ demon. Physical quantities after $N$ feedback cycles are plotted. (a) [(b)] Number density in sublattice $A$ ($B$) for $N=40$. (c) [(d)] Density of the $x$ component of spin in sublattice $A$ ($B$) for $N=30$. (e) [(f)] Density of the $y$ component of spin in sublattice $A$ ($B$) for $N=15$. (g) [(h)] Dynamics of the total $x$ ($y$) component of spin. The initial state is given by $\hat{\rho}_{\mathrm{i}}=\ket{j_{0},A,+}\bra{j_{0},A,+}$ | with $j_{0}=L/2$ (see text). The parameters of the model are the same as in Fig. 18.
@@ VII.5 Chern demon
In the previous subsections, we have focused on the models of quantum feedback control that are characterized by point-gap topology. In this subsection, we present an example of topological Maxwell’s demon that exhibits nontrivial line-gap topology. Our main idea is to simulate the Haldane model [101] of a Chern insulator by using feedback control. We consider a spin-$1/2$ particle on a honeycomb lattice and perform a projective measurement of the position of the particle so that the measurement does not disturb the spin state as in Sec. VII.3 and VII.4. The measurement operator is thus given by
@@
where $J_{1}\in\mathbb{R}$ and $J_{2}\in\mathbb{R}$ are the hopping amplitudes for nearest-neighbor and next-nearest-neighbor sites, respectively. Note that this Hamiltonian $\hat{H}$ does not depend on a measurement outcome. Then, we perform the second projective measurement described by the measurement operator (166) to get an outcome $m_{2}=(\vec{j}_{2},a_{2})$. The set of measurement outcomes $\{m_{1},m_{2}\}$ informs us of the initial and final positions of the particle travelling under the Hamiltonian $\hat{H}$. Using this information of measurement outcomes, we apply feedback control with the unitary operator $\hat{U}_{m_{1},m_{2}}^{\prime}=\exp[-i\phi_{m_{1},m_{2}}\hat{S}^{z}]$, where $\phi_{m_{1},m_{2}}\in\mathbb{R}$ and $\hat{S}^{z}$ is the magnetization obtained by replacing $i$ with $\vec{i}$ in Eq. (145).
@@
similarly to Eq. (156), and each block is given by
@@
where $\sigma^{\prime}\neq\sigma$, $p(\vec{j}_{2},a_{2}|\vec{j}_{1},a_{1}):=|\bra{\vec{j}_{2},a_{2},\sigma}e^{-i\hat{H}\tau}\ket{\vec{j}_{1},a_{1},\sigma}|^{2}$, and $\mathrm{sgn}[\sigma]$ is defined by $\mathrm{sgn}[\uparrow]=-\mathrm{sgn}[\downarrow]=+1$. Equation (170) indicates that one can imprint an arbitrary phase factor $e^{-i\phi_{m_{1},m_{2}}}$ on the transition amplitude $p(\vec{j}_{2},a_{2}|\vec{j}_{1},a_{1})$ by using feedback control. From a general perspective, this gives a significant advantage to realize topological phases with feedback control. In the present case, we set $\phi_{m_{1},m_{2}}$ to be nonzero if and only if $m_{1}$ and $m_{2}$ are nearest-neighbor sites, and we choose $\phi_{m_{1},m_{2}}=\pm\phi\in\mathbb{R}$, where the plus (minus) sign is taken if the hopping from $m_{1}$ to $m_{2}$ is in the counterclockwise (clockwise) direction. This choice of the phase factor resembles that in the Haldane model [101], thereby making the topologically trivial model (167) topological. Because of the phase factor, the block $\mathcal{E}_{\sigma\sigma^{\prime}}$ belongs to class A, while the block $\mathcal{E}_{\sigma\sigma}$ belongs to class AI due to the modular conjugation symmetry.
@@
In Fig. 20, we plot the dispersion relation of the blocks $\mathcal{E}_{\uparrow\uparrow}$ and $\mathcal{E}_{\uparrow\downarrow}$ of the Chern demon constructed above. As seen from the figure, the phase $\phi$ imprinted by the feedback control opens a gap in the eigenspectrum of $\mathcal{E}_{\uparrow\downarrow}$, and gapless edge modes that connect the bulk bands emerge under the OBC. This is analogous to the physics of the Haldane model. Thus, the quantum channel is characterized by nontrivial line-gap topology of $\mathcal{E}_{\uparrow\downarrow}$, which results in the emergence of the boundary-localized edge modes.
@@
Figure 20: Dispersion relations of the real part of eigenvalues of (a) the block $\mathcal{E}_{\uparrow\uparrow}$ and (b) the block $\mathcal{E}_{\uparrow\downarrow}$ of a quantum channel for the Chern demon. The parameters of the model are set to be $J_{1}=1,J_{2}=0.3,\tau=0.5$, and $\phi=\pi/2$. The unit of length is chosen to be the distance between nearest-neighbor sites. We employ the armchair edge in which one direction is periodic and the other direction is open with ten unit cells. The horizontal axis is the momentum along the edge. The eigenvalues of edge-localized states are colored in red in panel (b). The imaginary part of the eigenvalues is negligible.
@@
Here, a special attention should be paid on a physical difference between a Chern insulator and the Chern demon presented in this section. A Chern insulator is characterized by line-gap topology of a Hamiltonian, which is a generator of the dynamics of a system. Since the derivative of the dispersion relation with respect to momentum gives the group velocity, the gapless spectrum of edge states leads to chiral modes, which propagate unidirectionally along the edges. In contrast, the Chern demon is defined from line-gap topology of a quantum channel, which is a time-evolution operator itself rather than its generator. Thus, the physical meaning of the gapless edge-mode spectrum of the Chern demon is different from that of a Chern insulator. In particular, the boundary-localized modes of the Chern demon do not necessarily propagate along the edges of the system. Its physical implication merits further investigation.

### v3

@@ VII.3 QND demon
Here, we present a model of topological feedback control with strong unitary symmetry by following the construction in Sec. VI.3. We consider a spin-$1/2$ particle on a 1D lattice with two sublattices. A quantum state at site $j$ with sublattice $a=A,B$ and spin $\sigma=\uparrow,\downarrow$ is denoted by |$\ket{j,a,\sigma}$. We perform a projective position measurement on this particle and assume that this measurement does not disturb the spin state of the particle. Thus, a measurement outcome is specified only by the site index $j$ and the sublattice index $a$; therefore, the measurement operator is given by
@@
is the magnetization. Thus, the position measurement does not influence the spin degrees of freedom, and it serves as a QND measurement for $\hat{S}^{z}$ (see Sec. VI.3.1).
@@
where $J_{1},J_{2}$ are the hopping amplitudes, $V$ is the height of the potential, and $h_{z}$ ($h_{s}$) denotes the strength of a uniform (staggered) magnetic field. The feedback Hamiltonian conserves the magnetization, i.e., $[\hat{H}_{m},\hat{S}^{z}]=0$. The Hamiltonians (146) and (147) are generalized versions of the chiral Maxwell demon discussed in Sec. IV. The CPTP map for this feedback control is given by Eq. (5) with $\hat{K}_{m}=\hat{U}_{m}\hat{M}_{m}$ and $\hat{U}_{m}=\exp[-i\hat{H}_{m}\tau]$.
Since $[\hat{M}_{m},\hat{S}^{z}]=[\hat{H}_{m},\hat{S}^{z}]=0$, the CPTP map for this feedback control has strong symmetry (see Sec. VI.3). Hence, the CPTP map can be decomposed into four parts that act on subspaces with different symmetry eigenvalues. Specifically, we have
@@
In Fig. 16, we plot the eigenspectra of the quantum channel $\tilde{\mathcal{E}}$ under the PBC and the OBC together with those of ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$. The topological characterization of the quantum channel can be applied to each symmetry sector. The PBC eigenspectrum of ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$ with $\sigma=\sigma^{\prime}$ contains a single loop with winding number $-2$, while that of ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$ with $\sigma\neq\sigma^{\prime}$ contains two loops, each of which has winding number $-1$. Here, the uniform and staggered magnetic fields lift the degeneracy of the eigenspectra of ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$’s. Importantly, the eigenspectrum of ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$ with $\sigma\neq\sigma^{\prime}$ is no longer symmetric with respect to the real axis, implying that the modular conjugation symmetry is broken in these symmetry sectors. Thus, ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$ with $\sigma\neq\sigma^{\prime}$ offers a topologically nontrivial superoperator in class A.
Figure 16: Eigenspectrum of a quantum channel for feedback control with strong unitary symmetry. In each figure, blue curves and the blue dot at the origin constitute the eigenspectrum under the PBC in the infinite-size limit, and orange dots form the eigenspectrum under the OBC with the number of unit cells $L=40$. (a) Full eigenspectrum of the CPTP map $\tilde{\mathcal{E}}$. (b) Eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\uparrow}$, which is identical to that of $\tilde{\mathcal{E}}_{\downarrow\downarrow}$. (c) Eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\downarrow}$. (d) Eigenspectrum of $\tilde{\mathcal{E}}_{\downarrow\uparrow}$. The parameters of the model are set to $J_{1}=J_{2}=1$, $V=+\infty$, $h_{z}=0.5$, and $h_{s}=1$. The duration of feedback is $\tau=2$.
@@ VII.4 $\mathbb{Z}_{2}$ demon
As shown in Sec. VI.3.3, a quantum channel in class AII requires unitary symmetry and $C$ symmetry with $\eta_{C}=-1$ of the truncated Bloch matrix. To construct feedback control that fulfills these requirements, we consider a spin-$1/2$ particle on a 1D lattice with two sublattices as in Sec. VII.3. We first perform a spin-nondisturbing projective joint measurement of the position and the sublattice, where the measurement operator is given by Eq. (144). For a measurement outcome $m_{1}=(j_{1},a_{1})$, we perform feedback control with Hamiltonian $\hat{H}_{m_{1}}$, which has symmetry
@@
Example of such feedback Hamiltonians are given by
@@
as schematically illustrated in Fig. 17. For computational simplicity, here we assume that $\hat{H}_{j_{1},a}$ acts only on three unit cells labeled by $j_{1}-1,j_{1},j+1$; however, this assumption is not essential. After the unitary time evolution described by $\hat{U}_{j_{1},a_{1}}=\exp[-i\hat{H}_{j_{1},a_{1}}\tau]$, we perform the second projective joint measurement of the position and the sublattice with the same measurement operators (144) and suppose that we obtain an outcome $m_{2}=(j_{2},a_{2})$. Then, using the set of measurement outcomes $m_{1}$ and $m_{2}$, we perform additional feedback control with the following unitary operator:
$\hat{U}_{m_{1},m_{2}}^{\prime}=\begin{cases}\exp[-i\pi\hat{S}^{z}]&\mathrm{if}\ j_{1}>j_{2}\ \mathrm{and}\ a_{1}\neq a_{2}\\ \hat{I}&\mathrm{otherwise}.\end{cases}$ (155)
This additional feedback control can be realized by a magnetic field, and it imprints a phase depending on the spin degrees of freedom. This feedback control is essential for realizing $C$ symmetry of a quantum channel (see below).
Figure 17: Schematic illustration of the $\mathbb{Z}_{2}$ demon. We first perform a projective joint measurement of the position and the sublattice without disturbing the spin degrees of freedom. In panel (a), if a particle is detected at site $j_{1}$ and sublattice $A$, we perform feedback control by using the Hamiltonian $H_{j_{1},A}$. In panel (b), if a particle is detected at site $j_{1}$ and sublattice $B$, we perform feedback control by using the Hamiltonian $H_{j_{1},B}$. After unitary evolution under the feedback Hamiltonian $H_{j_{1},a_{1}}$ $(a_{1}=A,B)$, we perform the second projective measurement and additional feedback with a magnetic field [see Eq. (155)].
@@
The unitary operator with the feedback Hamiltonian $\hat{H}_{m_{1}}$ satisfies the following reciprocity relations because of the symmetry between $\hat{H}_{j_{1},A}$ and $\hat{H}_{j_{1},B}$:
@@
which follows from the symmetry (151) of the feedback Hamiltonian. As a result, conditions (111a)-(111c) are satisfied, and thus ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$ belongs to class AII for $\sigma\neq\sigma^{\prime}$ and class AI $+$ psH+ for $\sigma=\sigma^{\prime}$.
The topological feedback control in class AII exhibits the $\mathbb{Z}_{2}$ skin effect [72]. To see this case, we show the eigenspectrum of the quantum channel for the $\mathbb{Z}_{2}$ demon under the PBC and the OBC in Fig. 18. The block-diagonalized CPTP map $\tilde{\mathcal{E}}_{\uparrow\uparrow}$ does not possess $C$ symmetry with $\eta_{C}=-1$, and therefore, the Kramers degeneracy of eigenvalues is not ensured. Consequently, the loop structure in the eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\uparrow}$ shown in Fig. 18(a) collapses near the real axis due to real-complex transitions, and it does not support nontrivial topology. In contrast, the block $\tilde{\mathcal{E}}_{\uparrow\downarrow}$ has $C$ symmetry with $\eta_{C}=-1$, leading to the Kramers-degenerated eigenspectrum shown in Fig. 18(b). The point-gap topology of this spectrum is characterized by a $\mathbb{Z}_{2}$ topological invariant [72]
@@
Indeed, the eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\downarrow}$ under the OBC forms a line rather than a closed loop, which is a hallmark of the $\mathbb{Z}_{2}$ skin effect. Interestingly, the $\mathbb{Z}_{2}$ skin effect can occur only in ~$\tilde{\mathcal{E}}_{\sigma\sigma^{\prime}}$ with $\sigma\neq\sigma^{\prime}$ since ~$\tilde{\mathcal{E}}_{\sigma\sigma}$ has modular conjugation symmetry and is therefore characterized by the winding number rather than the $\mathbb{Z}_{2}$ invariant (see Sec. VII.2). Physically, the eigenspectra in Fig. 18 indicate that off-diagonal elements of the density matrix in terms of the $\hat{S}^{z}$ basis are accumulated on the edges of the system, whereas diagonal elements are delocalized over the system. Such distinct behavior is induced by the additional feedback in Eq. (155), which nontrivially acts only on the off-diagonal elements.
Figure 18: Eigenspectrum of a quantum channel for the $\mathbb{Z}_{2}$ demon. In each figure, blue curves and the blue dot at the origin constitute the eigenspectrum under the PBC in the infinite-size limit, and orange dots form the eigenspectrum under the OBC with the number of unit cells $L=20$. (a) Eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\uparrow}$. (b) Eigenspectrum of $\tilde{\mathcal{E}}_{\uparrow\downarrow}$. The parameters of the model are set to $J_{1}=0.8,J_{2}=1$, and $\lambda=0.5$. The duration of feedback is $\tau=1$.
Here, we remark on the boundary condition of the $\mathbb{Z}_{2}$ demon. To see the non-Hermitian skin effect in symmetry-protected topological feedback control, a quantum channel under the OBC must respect the protecting symmetry. We give details on how to implement a symmetry-preserving OBC in Appendix E.
@@
In Fig. 19, we plot the number density and the transverse spin components of a particle after $N$ repeated applications of feedback control with the $\mathbb{Z}_{2}$ demon. The numerical results clearly show that the transverse spin components are localized at the edges due to the skin effect, whereas the number density is delocalized over the system. Thus, the $\mathbb{Z}_{2}$ skin effect of the topological feedback control manifests itself only in the off-diagonal components of the density matrix. Note that the transverse spin components are not conserved due to the additional feedback in Eq. (155), and they vanish in the steady state of the quantum channel [see Figs. 19(g) and 19(h)]. Therefore, the nontrivial topology of the off-diagonal sectors appears only in the transient dynamics under repeated applications of feedback control.
Figure 19: Dynamical signatures of nontrivial topology of a $\mathbb{Z}_{2}$ demon. We show physical quantities after $N$ feedback cycles. (a) [(b)] Number density in sublattice $A$ ($B$) for $N=40$. (c) [(d)] Density of the $x$ component of spin in sublattice $A$ ($B$) for $N=30$. (e) [(f)] Density of the $y$ component of spin in sublattice $A$ ($B$) for $N=15$. (g) [(h)] Dynamics of the total $x$ ($y$) component of spin. The initial state is given by $\hat{\rho}_{\mathrm{i}}=\ket{j_{0},A,+}\bra{j_{0},A,+}$ | with $j_{0}=L/2$ (see text). The parameters of the model are the same as in Fig. 18.
@@ VII.5 Chern demon
In the previous subsections, we have focused on the models of quantum feedback control that are characterized by point-gap topology. In this subsection, we present an example of a topological Maxwell demon that exhibits nontrivial line-gap topology. Our main idea is to simulate the Haldane model [101] of a Chern insulator by using feedback control. We consider a spin-$1/2$ particle on a honeycomb lattice and perform a projective measurement of the position of the particle so that the measurement does not disturb the spin state as in Secs. VII.3 and VII.4. The measurement operator is thus given by
@@
where $J_{1}\in\mathbb{R}$ and $J_{2}\in\mathbb{R}$ are the hopping amplitudes for nearest-neighbor and next-nearest-neighbor sites, respectively. Note that Hamiltonian $\hat{H}$ does not depend on a measurement outcome. Next, we perform the second projective measurement described by the measurement operator (166) to obtain an outcome $m_{2}=(\vec{j}_{2},a_{2})$. The set of measurement outcomes $\{m_{1},m_{2}\}$ informs us of the initial and final positions of the particle traveling under Hamiltonian $\hat{H}$. Using this information of measurement outcomes, we apply feedback control with the unitary operator $\hat{U}_{m_{1},m_{2}}^{\prime}=\exp[-i\phi_{m_{1},m_{2}}\hat{S}^{z}]$, where $\phi_{m_{1},m_{2}}\in\mathbb{R}$ and $\hat{S}^{z}$ is the magnetization obtained by replacing $i$ with $\vec{i}$ in Eq. (145).
@@
similar to Eq. (156), and each block is given by
@@
where $\sigma^{\prime}\neq\sigma$, $p(\vec{j}_{2},a_{2}|\vec{j}_{1},a_{1}):=|\bra{\vec{j}_{2},a_{2},\sigma}e^{-i\hat{H}\tau}\ket{\vec{j}_{1},a_{1},\sigma}|^{2}$, and $\mathrm{sgn}[\sigma]$ is defined by $\mathrm{sgn}[\uparrow]=-\mathrm{sgn}[\downarrow]=+1$. Equation (170) indicates that one can imprint an arbitrary phase factor $e^{-i\phi_{m_{1},m_{2}}}$ on the transition amplitude $p(\vec{j}_{2},a_{2}|\vec{j}_{1},a_{1})$ by using feedback control. From a general perspective, this approach gives a significant advantage to realize topological phases with feedback control. In the present case, we set $\phi_{m_{1},m_{2}}$ to nonzero if and only if $m_{1}$ and $m_{2}$ are nearest-neighbor sites, and we choose $\phi_{m_{1},m_{2}}=\pm\phi\in\mathbb{R}$, where the plus (minus) sign is taken if the hopping from $m_{1}$ to $m_{2}$ is in the counterclockwise (clockwise) direction. This choice of the phase factor resembles that in the Haldane model [101], thereby making the topologically trivial model (167) topological. Because of the phase factor, block $\mathcal{E}_{\sigma\sigma^{\prime}}$ belongs to class A while block $\mathcal{E}_{\sigma\sigma}$ belongs to class AI due to modular conjugation symmetry.
@@
In Fig. 20, we plot the dispersion relation of blocks $\mathcal{E}_{\uparrow\uparrow}$ and $\mathcal{E}_{\uparrow\downarrow}$ of the Chern demon constructed above. As seen from the figure, the phase $\phi$ imprinted by feedback control opens a gap in the eigenspectrum of $\mathcal{E}_{\uparrow\downarrow}$, and gapless edge modes that connect the bulk bands emerge under the OBC. This feature is analogous to the physics of the Haldane model. Thus, the quantum channel is characterized by nontrivial line-gap topology of $\mathcal{E}_{\uparrow\downarrow}$, which results in the emergence of boundary-localized edge modes.
@@
Figure 20: Dispersion relations of the real part of eigenvalues of (a) block $\mathcal{E}_{\uparrow\uparrow}$ and (b) block $\mathcal{E}_{\uparrow\downarrow}$ of a quantum channel for the Chern demon. The parameters of the model are set to $J_{1}=1,J_{2}=0.3,\tau=0.5$, and $\phi=\pi/2$. The unit of length is chosen to be the distance between nearest-neighbor sites. We employ the armchair edge in which one direction is periodic and the other direction is open with ten unit cells. The horizontal axis is the momentum along the edge. The eigenvalues of edge-localized states are colored in red in panel (b). The imaginary part of the eigenvalues is negligible.
@@
Here, special attention should be given to the physical difference between a Chern insulator and the Chern demon presented in this section. A Chern insulator is characterized by line-gap topology of a Hamiltonian, which is a generator of the dynamics of a system. Since the derivative of the dispersion relation with respect to momentum gives the group velocity, the gapless spectrum of edge states leads to chiral modes, which propagate unidirectionally along the edges. In contrast, the Chern demon is defined by line-gap topology of a quantum channel, which is a time-evolution operator itself rather than its generator. Thus, the physical meaning of the gapless edge-mode spectrum of the Chern demon is different from that of a Chern insulator. In particular, the boundary-localized modes of the Chern demon do not necessarily propagate along the edges of the system. This physical implication merits further investigation.